OpenStudy (anonymous):

solve the system of equations by graphing. 3x + y = 9 4x + 5y = 1

OpenStudy (anonymous):

y=-3x+9 y=-4/5x+1/5 graph these two lines

OpenStudy (anonymous):

9-3x = (1-4x)/5 45-15x=1-4x 44=11x x=4 y=-3

OpenStudy (anonymous):

Chingo...why'd you post the same question twice? I answered your other one, lol